OpenStudy (anonymous):

2^x+1 = 64 find x? and how you got the answer

OpenStudy (anonymous):

(x-1)² = 64 x-1 = √64 = ± 8 x = 1 ± 8 x = -7, 9

OpenStudy (turingtest):

is your question\[2^{x+1}=64\]or\[2^x+1=64\]???

OpenStudy (turingtest):

@Kdwndr please read the question more carefully, you seem not to have understood the problem.

OpenStudy (anonymous):

\[For, 2^{x}+1=64\] \[\rightarrow2^{x}=64-1\] \[\rightarrow \log 2^{x}=\log 63\] \[\rightarrow xlog2=\log63\] \[\rightarrow x= (\log 63)/(\log2)\]

jhonyy9 (jhonyy9):

64 equal 2 on exponent 6 so than x+1=6 x=5